OpenStudy (anonymous):

You have a credit card with an existing balance of $1,569.34, a credit limit of $2,000, and an interest rate of 12.5% APR. You pay $150.00 a month. When will the balance be below the target debt ratio to improve your credit score?

OpenStudy (anonymous):

@ganeshie8 Can you help, please?

OpenStudy (amistre64):

if i knew what the target debt ratio was ....

OpenStudy (anonymous):

Well it said under 50% for me, but I just couldn't seem to figure it out?

OpenStudy (amistre64):

if you make no new purchase then we can determine a new balance from the old balance/payments made using a formula.

OpenStudy (amistre64):

Bn = Bk^n - P(1-k^n)/(1-k) where k is a compounding term: (1 + apr/12) and n is the number of months

OpenStudy (anonymous):

Uhm..Alright, I think.

OpenStudy (amistre64):

does 50% mean that we want a balance of 100 or less?

OpenStudy (amistre64):

1000 that is

OpenStudy (anonymous):

They want you to never have a debt ratio of more than 50%

OpenStudy (amistre64):

1000 = 1569.34(1+.125/12)^n - 150(1-(1+.125/12)^n)/(1-(1+.125/12)) and solve for n lets see if its less clutter to do a general solution then plug in the specifics

OpenStudy (anonymous):

I'm sorry, I'm really confused.

OpenStudy (amistre64):

working a formula to solve for variable is all we are doing. if you have your own formula to work with thats fine, i just tend to use the one im familiar with

OpenStudy (anonymous):

No, it's fine. I don't really know what to do with it -- I'm just really bad with numbers. My apologies.

OpenStudy (amistre64):

Bn = Bk^n - P(1-k^n)/(1-k) Bn (1-k) = Bk^n (1-k) - P(1-k^n) Bn (1-k) = Bk^n (1-k) - P +Pk^n Bn (1-k) = k^n [B(1-k) +P] - P Bn (1-k) +P = k^n [B(1-k) +P] [Bn (1-k) +P]/[B(1-k) +P] = k^n log both sides log[Bn (1-k) +P]/[B(1-k) +P] = n logk log[Bn (1-k) +P]/[B(1-k) +P] ------------------------- = n logk

OpenStudy (amistre64):

now plugging in our numbers should give us a solution to n, the number of months it will take (log(1000(-.125/12)+150)-log(1569.34(-.125/12)+150))/(log(1+.125/12)) http://www.wolframalpha.com/input/?i=%28log%281000%28-.125%2F12%29%2B150%29-log%281569.34%28-.125%2F12%29%2B150%29%29%2F%28log%281%2B.125%2F12%29%29 with any luck thats typed in correctly :) and gets us to about 4 months, little over

OpenStudy (amistre64):

we can work it out in a spread sheet to see how the payments and such are stacking up for us

OpenStudy (amistre64):

end balance = beg balance(interest) - payment

OpenStudy (anonymous):

Sorry! Openstudy crashed and I just now got it back.

OpenStudy (amistre64):

yeah, its been being a pain for months now i believe.

OpenStudy (anonymous):

I've always had trouble with it..I think I get what you did, at least enough. Thank you so much for your help.

OpenStudy (amistre64):

youre welcome, good luck :)